Fermi National Laboratory


Accelerator Update for October 10, 2005 - October 12, 2005

Monday October 10
The day shift began with Operations monitoring a Recycler stash and the Antiproton Source (Pbar) stack, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store, with NuMI off while its Horn ground fault repairs finish, and with Booster down due to LCW and vacuum repairs.

A Linac technician began replacing the Power Amplifier (PA) for a RF station (LRF2) at 8:35 AM.

Operations prepared Linac for access at 8:54 AM. Experts replaced a magnet and other equipment.

A Preaccelerator expert reported at 11:37 AM that the I- Source LCW leak was fixed.

Operations reported at 11:52 AM that pressure had been restored to the Booster LCW system. Operators and RF techs began searching for water leaks.

Linac techs reported at 12:41 PM that they were replacing the PA for LRF3.

At 2:02 PM, CDF reported that their solenoid dumped its helium due to a chassis problem. Unfortunately, it will take sixteen hours for them to refill and cool down the solenoid.

At 2:44 PM, Booster experts reported a vacuum leak on one of the ORBUMP magnets. Techs began repairs.

At 3:32 PM, Operations reported that deck three of the Pelletron had failed. There are six high voltage decks in the Pelletron. Experts tried to compensate for the missing deck, but couldn't. That means no electron cooling until the Pelletron gets repaired.

At 5:13 PM, a Main Injector (MI) RF station (MIRF1) tripped off. The station would reset, but wouldn't stay on. Operators left the station off.

At 6:09 PM, LRF3 tripped off and would not reset. An expert investigated and had the station back up by 7:38 PM.

At 8:17 PM, LRF3 tripped off for the first of two times in two minutes. It reset okay each time.

CDF came out of access at 8:40 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 10:36 PM.

Operations reported that the Booster RF was back up and running at 10:43 PM.

Tuesday October 11
The midnight shift began with Operations monitoring the stash and stack, and with all other downstream beam off due to Booster vacuum work.

Booster vacuum returned to normal operations and Operations had established beam to the Booster dump by 2:16 AM.

MiniBooNE began taking beam at 2:22 AM.

Pbar began stacking at 2:35 AM.

Operations put the TeV through a wet squeeze at 2:47 AM.

Operations established beam to the Meson MCenter experiment at 3:02 AM.

Operations put the TeV into shot setup at 4:18 AM.

Operations established combination store 4439 at 6:10 AM with an initial luminosity of 122.06E30.

Operations began preparing the Recycler at 6:56 AM to stash antiprotons.

At 8:06 AM, LRF3 tripped off. An expert began babysitting the station.

The Recycler stash ended and Pbar and MCenter had resumed normal operations by 8:20 AM.

Operations reported at 11:32 AM, that a problem with a Booster kicker (MKS04) had developed earlier this morning. Operations began preparing the Booster for an access.

Linac techs reported at 11:42 AM that they would change a power tube on LRF3 during the Booster shutdown.

At 11:52 AM, the Feynman Computer Center reported a power outage. Operations could find no evidence of a power glitch.

The Duty Electrician reported at 12:37 PM that a transformer trip had caused the Feynman Center power outage. They reset the transformer.

At 1:18 PM, Operations disabled stacking so a tech could work in an area behind the Pbar AP0 service building. There was no stacking anyway due to Booster work.

Booster experts and Operations began the Booster turn on at 3:06 PM.

Pbar resumed stacking at 3:25 PM.

MiniBooNE and MCenter resumed taking beam at 3:29 PM.

Operations terminated store 4439 at 6:04 PM.

Cryo system techs began work at 6:20 PM on the TeV sector E1 wet engine.

Operations prepared CDF and D0 for access into their collision halls at 6:20 PM.

CDF came out of access at 6:57 PM. D0 came out of access at 7:27 PM.

The Recycler began preparing to stash at 7:32 PM, but then never did the stash.

Operations put the TeV through a dry squeeze at 7:35 PM.

Pbar and MCenter resumed normal operations by 8:30 PM.

At 8:53 PM, the cryo techs reported that they found the E1 wet engine had a bad shaft that needed to be replaced.

The Recycler began preparing to stash at 8:53 PM. The first transfer of antiprotons was lost, but the rest of the stash was completed by 9:47 PM. Pbar resumed stacking.

A cryo tech reported that they had completed their work on the E1 wet engine shaft at 10:39 PM.

Wednesday October 12
The midnight shift began with Operations monitoring the stash, with Pbar stacking, with NuMI off, and with MiniBooNE and MCenter taking beam.

At 12:06 AM, LRF2 tripped off. An operator called a Linac expert for advice. Operators managed to get the station back on line by 1:17 AM.

Operations put the TeV through a dry squeeze at 1:30 AM.

Operations put the TeV into shot setup at 2:28 AM.

Operations established a Recycler only store 4442 at 4:14 AM with an initial luminosity of 110.14E30. MCenter resumed taking beam.

At 5:09 AM, LRF2 tripped off. It reset okay.

At 7:14 AM, LRF3 tripped off for the first of two times in thirty minutes. It reset both times.

The Plans for Wednesday
The plans for today are to allow an access into MI at 11 AM to work on a NuMI kicker. The work is expected to take one and half hours. The stash will have to be dumped for this access. D0 experts reported that they had replaced their faulty helium relief valve and they would power their solenoid this afternoon. NuMI experts reported that if things go good, they might start taking beam this evening.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/12/2005   email Fermilab